OpenStudy (anonymous):

\[\frac{ \sqrt{5\sqrt{5}} }{ \sqrt[8]{25} }\]

OpenStudy (anonymous):

\[\frac{ \sqrt{5\sqrt{5}} }{ \sqrt[8]{25} }\]